OpenStudy (eskelle):

Which cosine function has maximum of 2, a minimum of -2, and a period of 2pi/3?

OpenStudy (eskelle):

i tried to google this and it came up with weird stuff about scheduling. this is worded oddly and idk how to solve it

OpenStudy (mathmate):

Hints: A typical cosine function can be described as follows: \(\Large f(x)=Acos(B(x-C))+D\) A is the amplitude, equal to (maximum-minimum)/2, it acts as a vertical scale factor. B is relates to the period, and period equals \(2\pi/B\), works as a horizontal scale factor. C is horizontal shift. D is vertical shift. In the given case, C and D are zero, because there is no shift (translation). Normally cos(x) has a maximum of +1 and minimum of -1. A will amplify cos(x) to the given values (vertical scale factor). To find B, you would solve the equation: \(\Large \frac{2\pi}{B}=\frac{2\pi}{3}\) For more information and examples, read: http://www.regentsprep.org/regents/math/algtrig/att7/sinusoidal.htm

OpenStudy (eskelle):

okay, i still dont really understand though

Vocaloid (vocaloid):

well, going off of what mathmate said

Vocaloid (vocaloid):

our basic template is Acos(B(x−C))+D

Vocaloid (vocaloid):

the maximum/minimum of the basic cos function is 1 and -1, but in this case our minimum/max is -2 and 2 instead

Vocaloid (vocaloid):

that means, we need to multiply the entire function by 2

OpenStudy (eskelle):

what is the function

Vocaloid (vocaloid):

this Acos(B(x−C))+D

Vocaloid (vocaloid):

we go from Acos(B(x−C))+D to 2cos(B(x-C)) + D

OpenStudy (eskelle):

okay

Vocaloid (vocaloid):

C is our horizontal shift, and since we don't have a horizontal shift, C = 0

Vocaloid (vocaloid):

D is vertical shift, and we don't have one here either, so D = 0

Vocaloid (vocaloid):

that means we go from 2cos(B(x-C)) + D to 2cos(Bx) last step is to find B

Vocaloid (vocaloid):

our equation is 2pi/B = period

Vocaloid (vocaloid):

so since period = 2pi/3 2pi/B = 2pi/3 what does B equal?

OpenStudy (eskelle):

3

Vocaloid (vocaloid):

great

Vocaloid (vocaloid):

so our final function is 2cos(3x)

Vocaloid (vocaloid):

that's it

OpenStudy (eskelle):

oh okay

OpenStudy (eskelle):

thank you